Grade 12Physical Chemistry

calculate the osmotic presure of a solution obtain by mixing 100ml of solution of 1.5 %of urea and 100ml of solution of 3.42 %of cane sugar at 20°celsius?

Rajkaran singh
9 Years agoGrade 12

2 Answers

varun
ApprovedApproved Tutor Answer9 Years ago
HELLO RAJKARAN,
Osmotic pressure of solution 
100 ml of 1.5% urea
and 100ml 3.42% of cane sugar
Molar concentration in 200ml,we have 
solutes = 1.5/60 moles urea + 3.42/342 moles cane sugar
             =  0.025 + 0.01
              = 0.035 moles
                  => 0.035 X 1000
                      -----------------------
                             200
               => 0.175 M
Since cane sugar and urea are non ionic we will not multiply by the any factor
therefore,
pi = cRT 
    =  0.175 X 8.314 X 293
     =  426.335 Pa
THANK YOU
Aman Raj
7 Years ago
 
Osmotic pressure of solution 
100 ml of 1.5% urea
and 100ml 3.42% of cane sugar
Molar concentration in 200ml,we have 
solutes = 1.5/60 moles urea + 3.42/342 moles cane sugar
             =  0.025 + 0.01
              = 0.035 moles
                  = 0.035 X 1000/ 200
               = 0.175 M
Since cane sugar and urea are non ionic we will not multiply by the any factor
therefore,
pi = cRT 
 
= 0.175 × 0.0821 × 293
 
=4.209 atm